A copier was purchased by a company for $9,500. After 2 years it is estimated that the value of the copier will be $8,300. If the value in dollars V and the time the copier has been in use t are related by a linear equation, find the equation that relates V and t.
Answer by stanbon(75887)

A copier was purchased by a company for $9,500. After 2 years it is estimated that the value of the copier will be $8,300. If the value in dollars V and the time the copier has been in use t are related by a linear equation, find the equation that relates V and t.
---------
You have 2 points relating V and t: (0,9500) and (2,8300)
-----
slope = (8300-9500)/(2-0) = -1200/2 = -600
-----
intercept: 9500
------
Equation:
V = -600t + 9500
====================
cheers,
Stan H.
====================